OpenStudy (anonymous):

the perimeter of a triangle cannot be more thAN 46 FT. If one side of a triangle is twice as long as the second side and the third side is 19ft long. Find the longest possible values for the other two sides

OpenStudy (btaylor):

P = a + b + c if a = 19, we know that b=2c. So: 46 = 19 + b + c --> 27 = 2c+c 27 = 3c --> c = 9 since b=2c, b=2(9) = 18. 9 and 18 are your answers.

OpenStudy (anonymous):

awesome thank you taylor

OpenStudy (btaylor):

yw